OpenStudy (anonymous):

Can someone please help me with this problem? I am completely lost on how to solve this. 2x -1/3 + 3 < or = -4

OpenStudy (anonymous):

\[\frac{ 2x-1 }{ 3 } +3 \le -4\] That's how I meant, but I'm new here and didn't realize I could do this with the equation.

OpenStudy (anonymous):

\[(2x -1)\div3 <= -4 -3\]

OpenStudy (anonymous):

\[(2x-1) <= (-7)(3)\] \[(2x-1) <= -21\]

OpenStudy (anonymous):

= \[2x <= -21 + 1\]

OpenStudy (anonymous):

\[2x < -20\] \[x <= 20/2\]

OpenStudy (anonymous):

\[x<=10\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Anytime.